# Face-Human-Bench: A Comprehensive Benchmark of Face and Human Understanding for Multi-modal Assistants

### v2

@@ 3.1 Experimental Setup
Gemini-1.5-Pro 50.0 75.0 52.0 25.3 38.7 74.0 84.0 79.0 82.0 55.6 64.9 52.8 71.3 60.3
Claude-3.5-Sonnet 71.5 90.0 54.0 42.7 48.3 74.0 80.0 77.0 74.0 62.9 72.2 70.0 68.4 67.5
GPT-4V 73.0 78.0 38.0 71.3 54.7 68.0 84.0 76.0 83.0 65.7 72.9 66.4 73.7 69.3
GPT-4o 63.5 81.0 50.0 58.7 54.3 66.0 94.0 80.0 79.0 68.5 71.6 68.9 71.7 70.0
@@ 3.2 Main Results
Table 1 shows the performance of all evaluated MLLMs at different levels of abilities on the Human-Face-Bench (English)  [^4] For the results of the Chinese version, please refer to Appendix C.2. under the zero-shot setting. Overall scores range from 27.9% to 76.4%, demonstrating the effectiveness of the Face-Human-Bench in distinguishing the abilities of MLLMs in face and human understanding. We visualize the overall scores of MLLMs in Figure 2. Our findings can be summarized as follows.
@@
Overall Performance. (1) The top three performing open-source models in terms of the overall score are InternvL-Chat-v1.2-Plus, LLaVA-Next-34B, and InternVL-Chat-v1.5. These models’ LLMs have the largest number of parameters among all open-source models we evaluate. (2) Generally, open-source models within the same series tend to show improved performance with increasing parameter scale. However, there are exceptions; for instance, the 13B version of LLaVA-1.5 and LLaVA-Next perform slightly worse than their 7B counterparts. (3) Under the zero-shot setting, the best closed-source model, GPT-4o, does not surpass the performance of the top-performing open-source models. We believe this is because GPT-4o does not fully realize its potential under the zero-shot setting. The experiments in Section 3.5 confirm our hypothesis. (4) Newer models show significant improvements compared to earlier models. Among MLLMs with 7B parameters within LLM, the recently released LLaVA-OneVision-7B performs best. Impressively, LLaVA-OneVision-0.5B, with only 0.5B parameters within LLM, outperforms the earlier InstructBLIP-13B.
@@
L2 and L3 Performance [^5] For the visualization of L2 and L3 results, please refer to the Appendix C.1. (1) At L2 and L3, the best performance among open-source models is usually achieved by one of InternvL-Chat-v1.2-Plus, LLaVA-Next-34B, and InternVL-Chat-v1.5. Specifically, GLM-4V-9B achieves the best results in compound expression recognition (L3), facial expression recognition (L2), and action recognition (L2) and CogVLM2-19B-Chat achieves the best result in relative position understanding (L3). (2) At L2 and L3, the best performance among closed-source models is usually achieved by GPT-4o or GPT-4v. Notably, Gemini-1.5-Pro demonstrates outstanding face recognition ability (L2), achieving the best performance among all models with a score of 85.6%.
3.3 Correlation Between Abilities
In this section, we examine whether improving one ability in a model will enhance another by calculating the Pearson Correlation Coefficient between abilities at different levels, using the evaluation scores from Section 3.2. At L1, the correlation coefficient of face and human understanding is 0.94 and the correlation coefficient of perception and reasoning is 0.79, both indicating significant positive correlations, as shown in Figure 3(a) and Figure 3(b). We further investigate the correlations between L2 abilities, resulting in the correlation coefficient matrix shown in Figure 3(c). For clarity, we have drawn this as a lower triangular matrix. Our findings can be summarized as follows: (1) For the three face understanding abilities—facial attribute recognition, age estimation, and facial expression recognition—there are high positive correlations between each pair. (2) For the four human understanding abilities—human attribute recognition, action recognition, spatial relation understanding, and social relation understanding—there are high positive correlations between each pair. (3) For the three face understanding abilities and four human understanding abilities mentioned above, there are high positive correlations between each pair. (4) The two identity recognition tasks—face recognition and person re-identification—show a high positive correlation. (5) The correlation between face attack detection and any other ability is low. In Appendix C.3, we further present the correlations between L3 abilities.
@@ 3.4 Relative Position of Targets
Figure 4: (a) The versions used for the three face understanding abilities. (b) The versions used for human attribute recognition. (c) When MLLMs are evaluated with different versions, the wording of the questions varies slightly.
We investigate the impact of the relative position of targets on performance in four L3 abilities: facial attribute recognition, age estimation, basic expression recognition, and human attribute recognition. As shown in Figure 4, for the three face understanding abilities, we provide both the original and cropped versions, where only one person is included but the relative position varies. For human attribute recognition, we offer box-added and cropped versions. In the box-added version, multiple people are included, with the target to be discussed indicated by a red box. Figure 5 illustrates the performance differences between the two versions across various models. Our findings can be summarized as follows.
@@
We define the relative position sensitivity score (RPSS) as the sum of the absolute differences in scores between the two versions across the four tasks. This metric can serve as an effective reference for training MLLMs with more robust visual alignment for face and human understanding. We observe that InternLM-XComposer2-VL-7B, LLaVA-OneVision-7B, InternVL-Chat-v1.5, LLaVA-NeXT-34B, and InternVL-Chat-v1.2-Plus not only perform well in the four tasks but also exhibit low sensitivity scores. Among them, InternLM-XComposer2-VL-7B has the lowest sensitivity score of only 3.7%.666 For more models’ RPSS, please refer to the Appendix C.4.
@@ 3.5 CoT prompting
Table 2: Scores of the best open-source model, InternVL-Chat-v1.2-Plus, and the best closed-source model, GPT-4o, under different settings. ZS is short for Zero-Shot, H is short for Hints, VCoT is short for Vanilla CoT, 1TCoT is short for 1-stage Task-specific CoT. 2TCoT is short for 2-stage Task-specific CoT. Q is short for Question. O is short for Options. A is short for Answer. R is short for Relevant Analysis. The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively.
@@
In this section, we select InternVL-Chat-v1.2-Plus and GPT-4o to explore whether incorporating hints and Chain-of-Thought (CoT) instructions in the prompts can enhance the MLLMs’ performance. These two models have achieved the best overall performance in the main experiment among open-source models and closed-source models respectively. A hint involves tips on how to answer the question. For example, the hint for person re-identification is “if two people have significant differences in posture and their faces are relatively blurry, the main basis for determining whether they are the same person is their clothing characteristics.” CoT instructions, on the other hand, guide MLLMs to articulate the reasoning process leading to the answer. The vanilla CoT instruction simply requires the model to “analyze the question and options step by step”, whereas task-specific CoT instructions provide more tailored guidance based on the task. For example, for the deepfake detection task, the prompt might instruct the model to “analyze whether there are any artifacts in the facial image.” Following Multi-modal CoT (Zhang et al., 2024), we also conduct ablation experiments with both 1-stage and 2-stage frameworks. In the 1-stage framework, MLLMs are required to sequentially output the relevant analysis (rationale) and the answer in one round of dialogue. In the 2-stage framework, MLLMs first output the relevant analysis (rationale) in the first round and then provide the answer in the second round. Hints and task-specific CoT instructions for each L3 ability can be found in Appendix B.2.3.
Table 2 presents the performance of InternVL-Chat-v1.2-Plus and GPT-4o after incorporating hints and three different CoT settings. The results indicate that including hints and CoT instructions does not improve the performance of the open-source model; in fact, it may even cause a slight performance decline. By analyzing the outputs, we find that the open-source model does not provide rationales in its responses after adding CoT instructions to prompts. We believe this could be due to the model’s insufficient generalization capabilities, preventing it from understanding the CoT instructions. In contrast, the closed-source GPT-4o shows significant performance improvements. Adding hints leads to a 3.4% improvement compared to the zero-shot setting. Building upon this, vanilla CoT, 1-stage task-specific CoT, and 2-stage task-specific CoT further improve performance by 5.2%, 6.5%, and 5.7%, respectively. Ultimately, the combination of hints and 1-stage task-specific CoT instructions emerge as the best setting for overall performance.
In Figure 6, we further explore the main reasons for the performance improvements of GPT-4o in each ability at L2. Hints significantly improve performance in face attack detection, face recognition, and person re-identification, while CoT instructions significantly improve performance in facial attribute recognition, face attack detection, human attribute recognition, and action recognition. For the reasons behind the performance improvements in each ability at L3, please refer to Appendix C.5.
@@ 3.6 Specialist Models Significantly Outperforming MLLMs
In this section, we explore whether specialist models corresponding to 13 L3 abilities can be used to enhance MLLMs.  [^8] We explain the reasons for not conducting experiments on the remaining 5 L3 abilities in Appendix B.3.1. We directly test the performance of MLLMs using original datasets from the face and human community to facilitate comparison with specialist models. We design a set of prompt templates to transform the classification problems into multiple-choice problems and the regression problems (age estimation and crowd counting) into fill-in-the-blank problems.  [^9] For prompt templates, please refer to Appendix B.3.2. Specialist models are generally trained and tested on data from the same distribution. They can achieve high performance even if the test labels contain noise. However, the visual information learned by MLLMs and the original datasets used for testing may exhibit data distribution bias. To enable an effective comparison, we utilize early specialist models (which emerged after the widespread adoption of deep learning) as a reference to judge the performance of MLLMs on these tasks. [^10] For the early specialist models used for comparison, please refer to Appendix C.6.
We further define the relative performance score $S$ to normalize performances across different tasks: $S=(P_{m}-P_{r})/(P_{s}-P_{r})$, where $P_{m}$ is the performance of the MLLM. Here, we take the highest-performing model among InternVL-Chat-v1.2-Plus, LLaVA-Next-34B , and InternVL-Chat-v1.5 (the top three models in the main experiment). $P_{r}$ is the performance of random responses, and $P_{s}$ is the performance of the early specialist model. This metric typically ranges from 0 to 1, where a higher relative score indicates stronger abilities of MLLMs on the corresponding task. A relative score below 0 stands for even worse results than random responses, whereas a score above 1 indicates the performance surpassing the corresponding specialist models for reference. As shown in Table 3, there is no need for MLLMs to introduce specialist models to enhance the response quality when the abilities of age estimation, facial expression recognition, face anti-spoofing, action recognition, and person re-identification are needed. In contrast, for deepfake detection and crowd counting tasks, the MLLM significantly underperforms specialist models. Moreover, for face recognition, MLLMs can approach the specialist model under the basic scenario but indicate poor performance under more challenging scenarios, such as cross-pose, cross-age, similar-looking, and occluded. To sum up, we recommend incorporating the corresponding specialist models into multi-modal assistants for applications where deepfake detection, crowd counting, and accurate face recognition are required. Appendix F provides a demonstration of how to enhance multi-modal assistant responses with specialist models.
@@ 4 Related Work
Evaluation of MLLMs about Face and Human Understanding. Currently, there is no dedicated benchmark evaluating the face and human understanding abilities of MLLMs. Some efforts aim at comprehensively benchmarking MLLMs, containing some ability dimensions about face and human understanding. LAMM (Yin et al., 2023) evaluates 9 different 2D vision tasks using 11 existing public datasets. Among these, the facial classification task utilizes the CelebA (Liu et al., 2015) dataset to evaluate the accuracy of smile detection and hair attribute classification. MME (Fu et al., 2023) includes the celebrity recognition ability, requiring MLLMs to respond with Yes/No answers. SEED-Bench (Li et al., 2023a) includes the action recognition ability, where the inputs consist of multiple frames taken from a video, and MLLMs are required to choose the correct answer from four descriptions. MMBench (Liu et al., 2023c) includes the most extensive set of abilities related to faces and humans: celebrity recognition, action recognition, identity reasoning, and social relation, all of which are tested using multiple-choice problems. Considering the importance of faces and humans in multimedia, these evaluations are insufficient.
Face and Human Understanding. Face and human understanding is among the earliest research topics in artificial intelligence with successful applications. During the 2010s, the introduction of deep learning, particularly convolutional neural networks, significantly advanced face and human perception. In that era, numerous high-quality datasets were proposed for training and evaluating tasks of face attribute recognition (Liu et al., 2015), age estimation (Rothe et al., 2015; Escalera et al., 2015; Zhang et al., 2017), facial expression recognition (Barsoum et al., 2016; Li et al., 2017b; Mollahosseini et al., 2019), deepfake detection (Rössler et al., 2019; Dolhansky et al., 2019), face anti-spoofing (Liu et al., 2018; 2019), face recognition (Yi et al., 2014; Guo et al., 2016; Zheng et al., 2017; Deng et al., 2017; Zheng & Deng, 2018), human attribute recognition (Li et al., 2016; Liu et al., 2017), human-object interaction detection (Gupta & Malik, 2015; Xu et al., 2019), crowd counting (Zhang et al., 2016), social relationship recognition Sun et al. (2017); Li et al. (2017a) and person re-ideitification Li et al. (2014); Zheng et al. (2015). Entering the 2020s, a new paradigm emerged, which initially pre-trains a task-agnostic backbone and then based on this, trains a unified face or human model (Ci et al., 2023; Wang et al., 2023b; Qin et al., 2024) to simultaneously handle multiple face and human understanding tasks within a unified structure. In our evaluation, we observe that in certain tasks, MLLMs do not perform as well as specialist models. Utilizing these unified face or human models as the specialist models to help MLLMs can greatly facilitate deployment.

### v3

@@ 3.1 Experimental Setup
Gemini-1.5-Pro Gemini-1.5 50.0 75.0 52.0 25.3 38.7 74.0 84.0 79.0 82.0 55.6 64.9 52.8 71.3 60.3
Claude-3.5-Sonnet Claude-3.5 71.5 90.0 54.0 42.7 48.3 74.0 80.0 77.0 74.0 62.9 72.2 70.0 68.4 67.5
GPT-4V GPT-4V 73.0 78.0 38.0 71.3 54.7 68.0 84.0 76.0 83.0 65.7 72.9 66.4 73.7 69.3
GPT-4o GPT-4o 63.5 81.0 50.0 58.7 54.3 66.0 94.0 80.0 79.0 68.5 71.6 68.9 71.7 70.0
@@ 3.2 Main Results
Table˜1 shows the performance of all evaluated MLLMs at different levels of abilities on the Human-Face-Bench (English) [^4] For the results of the Chinese version, please refer to Section C.2. under the zero-shot setting. Overall scores range from 27.9% to 76.4%, demonstrating the effectiveness of the Face-Human-Bench in distinguishing the abilities of MLLMs in face and human understanding. We visualize the overall scores of MLLMs in Figure˜2. The findings can be summarized as follows: (1) The top 3 performing open-source models in terms of the overall score are InternvL-Chat-v1.2-Plus, LLaVA-Next-34B, and InternVL-Chat-v1.5. These models’ LLMs have the largest number of parameters among all open-source models we evaluate. (2) Generally, open-source models within the same series tend to show improved performance with increasing parameter scale. However, there are exceptions; for instance, the 13B version of LLaVA-1.5 and LLaVA-Next perform slightly worse than their 7B counterparts. (3) Under the zero-shot setting, the best closed-source model, GPT-4o, does not surpass the performance of the top-performing open-source models. We believe this is because GPT-4o does not fully realize its potential under the zero-shot setting. The experiments in Section˜3.5 confirm our hypothesis. (4) Newer models show significant improvements compared to earlier models. Among MLLMs with 7B parameters within LLM, LLaVA-OneVision-7B performs best. Impressively, LLaVA-OneVision-0.5B, with only 0.5B parameters within LLM, outperforms the earlier InstructBLIP-13B.
L2 and L3 Performance [^5] For the visualization of L2 and L3 results, please refer to the Section C.1. (1) At L2 and L3, the best performance among open-source models is usually achieved by one of InternvL-Chat-v1.2-Plus, LLaVA-Next-34B, and InternVL-Chat-v1.5. Specifically, GLM-4V-9B achieves the best results in compound expression recognition (L3), facial expression recognition (L2), and action recognition (L2) and CogVLM2-19B-Chat achieves the best result in relative position understanding (L3). (2) At L2 and L3, the best performance among closed-source models is usually achieved by GPT-4o or GPT-4V. Notably, Gemini-1.5-Pro demonstrates outstanding face recognition ability (L2), achieving the best performance among all models with a score of 85.6%.
3.3 Correlation Between Abilities
In this section, we examine whether improving one ability in a model will enhance another by calculating the Pearson Correlation Coefficient between abilities at different levels, using the evaluation scores from Section˜3.2.
@@ 3.3 Correlation Between Abilities
At L1, the correlation coefficient of face and human understanding is 0.94 and the correlation coefficient of perception and reasoning is 0.79, both indicating significant positive correlations, as shown in Figure˜3(a) and (b). We further investigate the correlations between L2 abilities, resulting in the correlation coefficient matrix shown in Figure˜3(c). For clarity, we have drawn this as a lower triangular matrix. Our findings can be summarized as follows: (1) For the three face understanding abilities—facial attribute recognition, age estimation, and facial expression recognition—there are high positive correlations between each pair. (2) For the four human understanding abilities—human attribute recognition, action recognition, spatial relation understanding, and social relation understanding—there are high positive correlations between each pair. (3) For the three face understanding abilities and four human understanding abilities mentioned above, there are high positive correlations between each pair. (4) The two identity recognition tasks—face recognition and person re-identification—show a high positive correlation. (5) The correlation between face attack detection and any other ability is low. In Section˜C.3, we further present the correlations between L3 abilities.
@@ 3.4 Relative Position of Targets
Figure 4: (a) The versions used for the three face understanding abilities. (b) The versions used for human attribute recognition.
We investigate the impact of the relative position of targets on performance in four L3 abilities: facial attribute recognition, age estimation, basic expression recognition, and human attribute recognition. As shown in Figure˜4, for the three face understanding abilities, we provide both the original and cropped versions, where only one person is included but the relative position varies. For human attribute recognition, we offer box-added and cropped versions. In the box-added version, multiple people are included, with the target to be discussed indicated by a red box. Figure˜5 illustrates the performance differences between the two versions across various models. Our findings can be summarized as follows.
@@
We define the relative position sensitivity score (RPSS) as the sum of the absolute differences in scores between the two versions across the four tasks. This metric can serve as an effective reference for training MLLMs with more robust visual alignment for face and human understanding. We observe that InternLM-XComposer2-VL-7B, LLaVA-OneVision-7B, InternVL-Chat-v1.5, LLaVA-NeXT-34B, and InternVL-Chat-v1.2-Plus not only perform well in the four tasks but also exhibit low sensitivity scores. Among them, InternLM-XComposer2-VL-7B has the lowest sensitivity score of only 3.7%. [^6] For more models’ RPSS, please refer to the Section C.4
@@ 3.5 CoT prompting
In this section, we select InternVL-Chat-v1.2-Plus and GPT-4o to explore whether incorporating hints and Chain-of-Thought (CoT) instructions in the prompts can enhance the MLLMs’ performance. These two models have achieved the best overall performance in the main experiment among open-source models and closed-source models respectively. A hint involves tips on how to answer the question. For example, the hint for person re-identification is “if two people have significant differences in posture and their faces are relatively blurry, the main basis for determining whether they are the same person is their clothing characteristics." CoT instructions, on the other hand, guide MLLMs in articulating the reasoning process that leads to the answer. The vanilla CoT instruction simply requires the model to “analyze the question and options step by step", whereas task-specific CoT instructions provide more tailored guidance based on the task. For example, for the deepfake detection task, the prompt might instruct the model to “analyze whether there are any artifacts in the facial image." Following Multi-modal CoT Multimodal-CoT , we also conduct ablation experiments with both 1-stage and 2-stage frameworks. In the 1-stage framework, MLLMs are required to sequentially output the relevant analysis (rationale) and the answer in one round of dialogue. In the 2-stage framework, MLLMs first output the relevant analysis (rationale) in the first round and then provide the answer in the second round. Hints and task-specific CoT instructions for each L3 ability can be found in Section˜B.2.3.
Table˜2 presents the performance of InternVL-Chat-v1.2-Plus and GPT-4o after incorporating hints and three different CoT settings. The results indicate that including hints and CoT instructions does not improve the performance of the open-source model; in fact, it may even cause a slight performance decline. By analyzing the outputs, we find that the open-source model does not provide rationales in its responses after adding CoT instructions to prompts. We believe this could be due to the model’s insufficient generalization capabilities, preventing it from understanding the CoT instructions. Specifically, the reasons may include the following aspects: (1) The scarcity of vision-language training samples aligned with the evaluation tasks of face and human understanding. (2) The insufficient exposure to samples of the complex reasoning paradigm during the training process. (3) The inadequate model capacity, which stems from substantially smaller parameter sizes compared to those of closed-source MLLMs. In contrast, the closed-source GPT-4o shows significant performance improvements. Adding hints leads to a 3.4% improvement compared to the zero-shot setting. Building upon this, vanilla CoT, 1-stage task-specific CoT, and 2-stage task-specific CoT further improve performance by 5.2%, 6.5%, and 5.7%, respectively. Ultimately, the combination of hints and 1-stage task-specific CoT instructions emerge as the best setting for overall performance.
In Figure˜6, we further explore the main reasons for the performance improvements of GPT-4o in each ability at L2. Hints significantly improve performance in face attack detection, face recognition, and person re-identification, while CoT instructions significantly improve performance in facial attribute recognition, face attack detection, human attribute recognition, and action recognition. For the reasons behind the performance improvements in each ability at L3, please refer to Section˜C.5.
Table 2: Scores of InternVL-Chat-v1.2-Plus and GPT-4o under different settings. ZS is short for Zero-Shot, H is short for Hints, VCoT is short for Vanilla CoT, 1TCoT is short for 1-stage Task-specific CoT. 2TCoT is short for 2-stage Task-specific CoT. Q is short for Question. O is short for Options. A is short for Answer. R is short for Relevant Analysis. The highest scores for open-source and closed-source MLLMs are marked in blue and green respectively.
@@ 3.6 Specialist Models Significantly Outperforming MLLMs
In this section, we explore whether specialist models can significantly outperform MLLMs for the evaluation of 13 L3 abilities. [^7] We explain the reasons for not conducting experiments on the remaining 5 L3 abilities in Section B.3.1 We directly test the performance of MLLMs using original datasets from the face and human community to facilitate comparison with specialist models. We design a set of prompt templates to transform the classification problems into multiple-choice problems and the regression problems (age estimation and crowd counting) into fill-in-the-blank problems. [^8] For prompt templates, please refer to Section B.3.2 Specialist models are generally trained and tested on data from the same distribution. They can achieve high performance even if the test labels contain noise. However, the visual information learned by MLLMs and the original datasets used for testing may exhibit data distribution bias. To enable an effective comparison, we utilize early specialist models (which emerged after the widespread adoption of deep learning) as a reference to judge the performance of MLLMs on these tasks. [^9] For the early specialist models used, refer to Section C.6
We further define the relative performance score $S$ to normalize performances across different tasks: $S=(P_{m}-P_{r})/(P_{s}-P_{r})$, where $P_{m}$ is the performance of the MLLM. Here, we take the highest-performing model among InternVL-Chat-v1.2-Plus, LLaVA-Next-34B , and InternVL-Chat-v1.5 (the top three models in the main experiment). $P_{r}$ is the performance of random responses, and $P_{s}$ is the performance of the early specialist model. This metric typically ranges from 0 to 1, where a higher relative score indicates stronger abilities of MLLMs on the corresponding task. A relative score below 0 stands for even worse results than random responses, whereas a score above 1 indicates the performance surpassing the corresponding specialist models for reference.
As shown in Table˜3, MLLMs perform well in age estimation, facial expression recognition, face anti-spoofing, action recognition, and person re-identification, eliminating the need to introduce specialist models to enhance response quality. In contrast, for deepfake detection and crowd counting tasks, the MLLM significantly underperforms specialist models. Moreover, for face recognition, MLLMs can approach the specialist model under the basic scenario but indicate poor performance under more challenging scenarios, such as cross-pose, cross-age, similar-looking, and occluded. We recommend incorporating the corresponding specialist models into multi-modal assistants for applications where deepfake detection, crowd counting, and accurate face recognition are required. Appendix˜F provides a demonstration of how to enhance multi-modal assistant responses with specialist models.
@@ 4 Related Work
Evaluation of MLLMs for Face and Human Understanding. Currently, there is no dedicated benchmark evaluating the face and human understanding abilities of MLLMs. Some efforts aim at comprehensively benchmarking MLLMs, containing some ability dimensions about face and human understanding. LAMM LAMM evaluates 9 different 2D vision tasks using 11 existing public datasets. Among these, the facial classification task utilizes the CelebA CelebA dataset to evaluate the accuracy of smile detection and hair attribute classification. MME MME includes the celebrity recognition ability, requiring MLLMs to respond with Yes/No answers. SEED-Bench SEED-Bench includes the action recognition ability, where the inputs consist of multiple frames taken from a video, and MLLMs are required to choose the correct answer from four descriptions. MMBench MMbench includes the most extensive set of abilities related to faces and humans: celebrity recognition, action recognition, identity reasoning, and social relation, all of which are tested using multiple-choice problems. Considering the importance of faces and humans in multimedia, these evaluations are insufficient.
Face and Human Understanding. Face and human understanding is among the earliest research topics in artificial intelligence with successful applications. Numerous high-quality datasets were proposed for training and evaluating tasks of face attribute recognition CelebA , age estimation IMDB+WIKI ; CLAP2015 ; UTKFace , facial expression recognition FERPlus ; DLP-CNN-RAF-DB ; AffectNet , deepfake detection FFpp ; DFDC , face anti-spoofing SiW ; SiW-M , face recognition CASIA-WebFace ; MS-Celeb-1M ; CALFW ; SLLFW ; CPLFW , human attribute recognition WIDER ; PA-100K , human-object interaction detection V-COCO ; HICO-DET , crowd counting ShTech , social relationship recognition PIPA ; PISC and person re-ideitification CUHK03 ; Market-1501 . Entering the 2020s, a new paradigm emerged, which initially pre-trains a task-agnostic backbone and then based on this, trains a unified face or human model UniHCP ; Hulk ; Faceptor to simultaneously handle multiple face and human understanding tasks within a unified structure. In our evaluation, we observe that in certain tasks, MLLMs do not perform as well as specialist models. Utilizing these unified face or human models as specialist models to help MLLMs can greatly enhance responses.